a purse contains 4.65 cents in dimes and quaters. there are 13 more quaters than dimes. find the number of dimes and quaters
Answer by jorel1380(3719)

10d+25q=465
q=d+13
10d+25(d+13)=465
35d+325=465
35d=140
d=4
q=17
The purse has 4 dimes and 17 quarters..